# Fast Training Dataset Attribution via In-Context Learning

### v1

@@ 4 Preliminary Experiments
Tables 1 and 3 show the attribution results obtained by SCM and CMF algorithms, respectively. Both algorithms successfully identify the BoolQ dataset as the most influential dataset. This is because BoolQ context is more directly related to the questions. Chemistry, Natural Science, History and Archaeology, Biology, and Law have lower $\phi_{k}$ values, showing that while they contribute to the context, their impact is less significant compared to BoolQ. Note that in CMF, we need to calculate $\frac{\pi_{\text{BoolQ}}}{1-\pi_{\text{Base}}}$ to directly compare it with $\phi_{\text{BoolQ}}$ estimated by SCM.
@@
Table 1: Shapley Values ($\phi_{k}$) using SCM Algorithm
Comparing the results of SCM and CMF, we see that CMF places higher weight on BoolQ. With all LLMs, the quantity $\frac{\pi_{\text{BoolQ}}}{1-\pi_{\text{Base}}}$ is larger than $\phi_{\text{BoolQ}}$ obtained by SCM. We attribute this to the noise in RAG when we pull context from multiple sources. Moreover, CMF provides the contributions of the base LLM $\pi_{\text{Base}}$. We can see that GPT-4 has the highest contribution from the base language model. This observation is inline with the observation that GPT-4 has the highest accuracy without any context, as reported in Table 5 in the appendix.
We also conducted an evaluation using Trak (Park et al., 2023) as a popular baseline. Trak provides a different perspective on dataset attribution by scoring the impact of training data on model predictions. The Trak scores for Bloomz and Phi-3-mini across the same datasets are shown in Table 2 . Our results align well with Trak, and improve upon it, reinforcing the robustness of our methods.
Model BoolQ Chemist Natural Sci History Biology Law
Phi-3 0.58 -0.08 0.20 0.18 -0.05 0.07
Bloomz 0.61 -0.10 0.22 0.19 -0.06 0.09
By comparing our results with Trak, we observe that both methods identify BoolQ as the most significant contributor. However, our CMF approach provides a more detailed and accurate attribution of dataset contributions, particularly in quantifying the base model’s influence and managing the noise inherent in RAG systems.
Runtime Comparison
The CMF algorithm is faster than the SCM algorithm as it requires fewer queries with shorter context sizes. Utilizing an AWS EC2 G6 instance (g6.16xlarge), the total runtime for CMF, involving 7 runs, ranges from 77 to 94 minutes for all LLMs. In contrast, the SCM method, which requires $2^{5}$ runs, results in a total runtime of 352 to 384 minutes. Both algorithms’ runtimes are dominated by the RAG search time. This substantial reduction in runtime demonstrates the efficiency of the CMF method, making it more suitable for scenarios demanding both accuracy and computational efficiency.
Deep Dive into RAG Noise Effect
We compute the mean similarities and residuals for each model as shown in Table 4. The high residual for Bloomz (-0.32) indicates that BoolQ contributes significantly to the output. When the BoolQ context is added, the model’s performance changes markedly, demonstrating the influence of this dataset.
The low residual for GPT-4 (-0.03) suggests that GPT-4 has been pre-exposed to similar data. The minimal change in performance with the addition of BoolQ context indicates that GPT-4 already possesses substantial knowledge from similar datasets.
@@
Table 3: $\bm{\pi}$ and $\frac{\pi_{\text{BoolQ}}}{1-\pi_{\text{Base}}}$ values using CMF algorithm
The positive residual for Mistral 7B (0.06) implies that Mistral 7B relies on the added context. This reliance on context suggests that the model benefits greatly from the additional information provided by BoolQ, indicating that it has not been exposed to similar data during training. However, this also introduces RAG noise, as the model’s output can be significantly influenced by the context.
The small residual for Phi-3 (0.03) shows partial exposure to similar data during training. The model benefits from the added BoolQ context but to a lesser extent, suggesting that while it improves with additional context, it already has some degree of relevant knowledge.
Metric Bloomz GPT-4 Mistral 7B Phi-3
$s_{S}$ 0.86 (0.03) 0.76 (0.02) 0.73 (0.03) 0.60 (0.03)
$s_{S\cup\{D_{k}\}}$ 0.54 (0.02) 0.72 (0.02) 0.69 (0.03) 0.63 (0.02)
$r_{k,S}$0.32 (0.02) -0.03 (0.02) 0.06 (0.02) 0.03 (0.02)
@@
Table 4: Mean similarities $s_{S}$ and residuals $r_{k,S}$ for different LLMs (with standard deviations)
Context Bloomz GPT-4 Mistral 7B Phi-3
No Context 0.43 0.73 0.68 0.70
BoolQ as RAG 0.74 0.87 0.85 0.82
Five Datasets Only 0.35 0.64 0.60 0.45
All Datasets + BoolQ 0.73 0.84 0.82 0.83
Table 5: Accuracy of LLMs with Different Contexts
The baseline setting (No Context) reveals inherent differences in the LLMs’ capabilities. GPT-4 has the highest baseline accuracy at 0.73, followed by Phi-3 at 0.70 and Mistral 7B at 0.68, indicating robust pre-training for these models. Bloomz shows lower accuracy at 0.43, highlighting its dependency on contextual data.
When the BoolQ context is provided using RAG, all models show significant accuracy improvements, with GPT-4 reaching 0.87, Mistral 7B at 0.85, and Phi-3 at 0.82. Bloomz also improves to 0.74, though it remains lower than the others. Providing context from five datasets (excluding BoolQ) leads to accuracy drops for all models, with Bloomz at 0.35, GPT-4 at 0.64, Mistral 7B at 0.60, and Phi-3 at 0.45. This indicates that less relevant data is less effective for understanding BoolQ queries.
Combining all datasets with BoolQ context results in slight decreases in accuracy for GPT-4 (0.84) and Mistral 7B (0.82), suggesting that additional datasets introduce noise. Bloomz and Phi-3 show minimal changes, indicating that extra data does not significantly impact their performance once BoolQ context is included. These results emphasize the importance of relevant contextual information in enhancing LLM performance, with GPT-4 consistently outperforming other models due to its extensive training.
@@ 5 Conclusion and Discussion
Our results show that both of our proposed algorithms successfully attribute the output of LLMs to the BoolQ dataset (as the proxy for related knowledge). Comparing LLMs GPT-4 showed minimal change in the similarities when BoolQ context was added, suggesting prior exposure to similar data, while Bloomz exhibited a high residual, indicating substantial influence from BoolQ. The CMF algorithm provides further insights by quantifying the contributions of the base LLM. Comparing our two methods, we conclude that CMF is computationally less expensive and more robust to the RAG noise.
In this paper, we considered several datasets as proxies for the datasets used for training of LLMs. For future work, we will try to avoid this approximate method and train our LLMs on specific datasets and test our algorithms with ground truth contributing datasets obtained by retraining.
Acknowledgment
We would like to thank Layne Price and Bilal Fadlallah for their helpful comments.
Appendix A Appendix
Appendix B Synthetic Experiments with Matrix Factorization
To understand the impact of regularization terms in Eq. (7), we perform a simple synthetic experiment. We create a matrix of random $p(y|q)$ (uniformly between 0 and 1) for $m=5$ queries and decompose it to identify the impact of 3 sources. We vary the regularization terms $\lambda_{\pi}=\lambda_{P_{K}}\in[0,10^{-4},10^{-3},10^{-2},10^{-1},1]$. We repeat the experiments with randomly initialized $bm{\pi}$ and $P_{K}$ for 100 times and report the variations in the solutions for $\bm{\pi}$. The variation is measured as the sum of the eigenvalues of the covariance matrix of the solutions (i.e., nuclear norm). Figure 1 recommends choosing the regularization coefficients larger than $0.1$ to have stable solutions.
Figure 1: The Logarithm of the variations in the solutions for $\bm{\pi}$. The variation is measured as the sum of the eigenvalues of the covariance matrix of the solutions (i.e., nuclear norm). The shaded area shows the [5, 95] percentile.
Appendix C Algorithms
Algorithm 1 Shapley Context Method (SCM)
0:  An instruction-tuned LLM $M$ that generates outputs $y$ for each query $q$ and context $c$.
0:  A set of queries $Q=\{q_{1},\ldots,q_{m}\}$.
0:  A set of datasets that we need to compute their contributions $D_{1},\ldots,D_{K}$.
1:  for $q\in Q$ do
2:     Compute output without context: $y=M(q)$.
3:     for $S\subset\{D_{1},\ldots,D_{K}\}\setminus\{D_{k}\}$ do
4:        Use RAG to create contexts $c_{S}$ and $c_{S\cup\{D_{k}\}}$ from the datasets in $S$ and $D_{k}$.
5:        Compute the output with the context $y|c_{S}=M(q|c_{S})$.
6:        Compute the output with the context including $D_{k}$: $y|c_{S\cup\{D_{k}\}}=M(q|c_{S\cup\{D_{k}\}})$.
7:        Compute the similarities $s_{S}$ and $s_{S\cup\{k\}}$ using Eq. (2).
8:        Compute the residual gain $r_{k,S}=s_{S\cup\{k\}}-s_{S}$.
9:     end for
10:     Use Eq. (3) and $r_{k,S}$ to compute $\phi_{k}(q)$.
11:  end for
12:  Return average $\phi_{k}$ over $m$ queries.
Algorithm 2 Context Mixture Factorization (CMF)
0:  An instruction-tuned LLM $M$ that generates outputs $y$ for each query $q$ and context $c$.
0:  A set of queries $Q=\{q_{1},\ldots,q_{m}\}$.
0:  A set of datasets that we need to compute their contributions $D_{1},\ldots,D_{K}$.
1:  for $q\in Q$ do
2:     Compute output without context: $p[y|q,c_{0}]=M(q)$.
3:     for $k=1,\ldots,K$ do
4:        Use RAG to create context $c_{k}$ from the dataset $D_{k}$.
5:        Compute the output with the context $p[y|q,c_{k}]=M(q|c_{k})$.
6:     end for
7:  end for
8:  Build matrix $P$, where $P_{k,j}=p[y|q_{j},c_{k}]$.
9:  Solve Eq. (7) via alternating least squares and to compute $\widehat{\bm{\pi}}$.
10:  Return the contribution vector $\widehat{\bm{\pi}}$.
Appendix D Implementation Details
D.1 Prompt Engineering
For simplicity of evaluation and without loss of generality, we used BoolQ (Clark et al., 2019) Q&A dataset, where the answers are binary Yes/No. To instruct the LLMs to provide direct boolean responses, we used prompt engineering. Initially, we tested various prompts without explicitly instructing the model to answer with ”Yes” or ”No.” Diverse examples used in this process are provided in Appendix D.3. Through iterative testing, we found that responses improved when the model was explicitly instructed to provide a boolean answer. This led to our final prompt:
Prompt: ”Given the context below, answer the question that follows with only ’Yes’, ’No’, or ’I don’t know’ if the context is insufficient.
{question}? The answer to this question is”
While this final prompt worked well for GPT-4, Bloomz, and Mistral 7B, generating straightforward ”Yes,” ”No,” or ”I don’t know” responses, it was harder to instruct Phi-3-mini. Even with the final prompt, Phi-3-mini often generated more text than just a simple boolean response.
D.2 Using RAG
Given the limitations of LLM context windows, fitting entire datasets directly into the context is impractical. To address this, we utilized Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) to enhance the context by retrieving relevant documents from databases before generating responses. The process involves splitting the documents into semantically relevant chunks using the RecursiveCharacterTextSplitter from the Hugging Face Transformers library, computing embeddings for all chunks with a model like thenlper/gte-small, and storing these embeddings in a vector database using FAISS (Facebook AI Similarity Search). When a question is posed, it is embedded and a similarity search is performed against the vector database to find the closest matching documents. These retrieved documents are then provided as context to the LLMs along with the original question, allowing the LLMs to generate responses augmented with additional context.
We used a chunk size of 512 and a top-k value of 3, ensuring the context was trimmed to 2000 characters for conciseness.
D.3 Prompts

### v2

@@ 4.1 Results and Analysis
Detailed Analysis of Attribution Coefficients Tables 4 and 5 show the attribution results obtained by the SCM and CMF algorithms for the BoolQ data set. Both algorithms successfully identify the BoolQ dataset as the most influential dataset. This is because the BoolQ context is more directly related to the questions. Chemistry, Natural Science, History and Archaeology, Biology, and Law have lower $\phi_{k}$ values, showing that while they contribute to the context, their impact is less significant compared to BoolQ. Note that in CMF, we need to calculate $\frac{\pi_{\text{BoolQ}}}{1-\pi_{\text{Base}}}$ to directly compare it with $\phi_{\text{BoolQ}}$ estimated by SCM. This shows that CMF assigns higher attribution values than SCM due to its robustness in accounting for noise in retrieval-augmented generation (RAG) systems.
Table 4: Shapley Values ($\phi_{k}$) using SCM Algorithm.
@@
Table 5: $\bm{\pi}$ and $\frac{\pi_{\text{BoolQ}}}{1-\pi_{\text{Base}}}$ values using the CMF algorithm.
@@
Dataset Phi-3 Bloomz
BoolQ 0.58 0.61
Chemistry -0.08 -0.10
Natural Science 0.20 0.22
History 0.18 0.19
Biology -0.05 -0.06
Law 0.07 0.09
Across the three datasets (see Tables 1013 in Appendix D), CMF consistently assigns higher attribution values than SCM. This is because CMF explicitly accounts for background model contributions ($\pi_{\text{Base}}$) and dataset-specific contributions ($\pi_{k}$), making it more robust to noise and improving attribution granularity. The difference between CMF and SCM is most pronounced for datasets with strong alignment with the pre-training data of the model, such as BoolQ, where CMF captures a clearer and stronger attribution signal. Since BoolQ contains questions and contexts similar to those likely encountered during model training, CMF detects these relationships with greater sensitivity.
For FakeQ and Olympic2024, both unseen during pre-training, the gap between CMF and SCM narrows. This is expected as neither data set has direct overlap with pre-training data, leading to lower attribution values across the board. However, CMF still assigns slightly higher attributions compared to SCM, particularly for FakeQ. This suggests that while FakeQ is designed to be unseen, it retains enough linguistic patterns and contextual structures resembling BoolQ for CMF to register a weak but measurable connection. In contrast, Olympic2024 shows the lowest attribution values, reflecting its novel and domain-specific nature. This trend underscores the ability of the CMF to differentiate datasets not only based on direct exposure but also through latent associations in linguistic or contextual patterns, making it a more reliable metric for evaluating dataset contributions.
These findings validate the superior sensitivity of CMF in identifying relevant data influences, even for datasets with no explicit pre-training overlap. At the same time, they illustrate that both methods converge to lower attribution values when applied to data sets entirely outside of the prior knowledge of the model, confirming the robustness of SCM and CMF in distinguishing between the data sets seen and the novel ones.
(a) SCM
(b) CMF
Figure 1: (left) SCM Attribution Values vs. Learning Rate for Olympic2024 Dataset: Attribution values increase with fine-tuning. (right) CMF Attribution Values vs. Learning Rate for Olympic2024 Dataset: CMF shows higher attribution values, reflecting its robustness during fine-tuning.
For BoolQ, Trak identified it as the most influential dataset, which aligns well with our methods. However, our CMF approach provides a more detailed and accurate attribution of dataset contributions, particularly in quantifying the base model’s influence and managing the noise inherent in retrieval-augmented generation (RAG) systems. CMF consistently assigns higher attribution values compared to SCM and Trak, reflecting its robustness in capturing the alignment between BoolQ and the model training data.
For FakeQ, Trak shows lower attribution scores compared to BoolQ, as expected for an unseen dataset. CMF again outperforms SCM and Trak by effectively taking advantage of the syntactic similarity of FakeQ to BoolQ, capturing its partial alignment with training data. SCM also performs reasonably well, but is less sensitive to subtle contributions, and Trak provides scores comparable to SCM, though it lacks the granularity CMF offers.
For Olympic2024, being entirely novel and unrelated to training data, all methods report significantly lower attribution values. CMF continues to demonstrate superior performance by reflecting even minimal data-set contributions while maintaining a clear distinction between seen and unseen data. SCM and Trak exhibit a smaller gap between Olympic2024 and FakeQ, indicating their limited ability to fully capture the novelty of datasets.
Case Study: Evaluation of Unlearning Methods
We applied attribution metrics to assess the effectiveness of machine unlearning methods. As shown in Table 7, three unlearning methods: Gradient Ascent (Golatkar et al., 2020; Liu et al., 2024), Fine-tuning with Random Labels (Golatkar et al., 2020), and Unlearning with Adversarial Samples (Cha et al., 2024) were evaluated based on their ability to reduce the influence of the Olympic2024 dataset on Bloomz and Phi-3 models.
Table 7: SCM and CMF Attribution Scores for the Olympic2024 Dataset.
Before Finetuning After Finetuning After Unlearning
Unlearning Method LLM SCM CMF SCM CMF SCM CMF
Gradient Ascent Bloomz 0.08 0.16 0.75 0.85 0.25 0.34
Phi-3 0.07 0.10 0.72 0.82 0.28 0.37
Fine-tuning Bloomz 0.08 0.16 0.75 0.85 0.22 0.31
with Random Labels Phi-3 0.07 0.10 0.72 0.82 0.24 0.33
Unlearning with Bloomz 0.08 0.16 0.75 0.85 0.30 0.41
Adversarial Samples Phi-3 0.07 0.10 0.72 0.82 0.33 0.42
The results indicate that Unlearning with Adversarial Samples consistently outperforms the other methods, achieving the highest reduction in attribution values for both SCM and CMF metrics. The ability of this method to target specific data points for unlearning is reflected in the reduced values for $\frac{\pi_{\text{Olympic}}}{1-\pi_{\text{Base}}}$. In contrast, Gradient Ascent and Fine-tuning with Random Labels achieve moderate reductions, with Gradient Ascent slightly outperforming Random Labels in most cases.
Runtime Comparison
The CMF algorithm is faster than the SCM algorithm as it requires fewer queries with shorter context sizes. Using an AWS EC2 G6 instance (g6.16xlarge), the total runtime for CMF, involving 7 runs, ranges from 77 to 94 minutes for all LLMs. In contrast, the SCM method, which requires $2^{5}$ runs, results in a total runtime of 352 to 384 minutes. The runtimes of both algorithms are dominated by the RAG search time. This substantial reduction in run-time demonstrates the efficiency of the CMF method, making it more suitable for scenarios demanding both accuracy and computational efficiency.
Validation of Attribution Metrics through Fine-Tuning
To establish the reliability of the attribution metrics, we performed a fine-tuning experiment on the Olympic2024 dataset. Fine-tuning was performed incrementally, with the learning rate increasing progressively. Figures 1a and 1b demonstrate that the attribution values increase monotonically with fine-tuning, confirming that the metrics effectively capture the growing dependence of LLM on the data set. For both SCM and CMF, the Olympic2024 attribution values increase with fine-tuning, but CMF shows higher attribution values due to its ability to handle noise more robustly. These results validate the ability of the proposed metrics to quantify the influence of fine-tuned data on LLMs.
The incremental increase in attribution values also reflects a nonlinear growth pattern, as the effects of fine-tuning diminish at higher learning rates, resulting in a plateau. This plateau effect is more prominent in SCM as it lacks the noise handling capabilities of CMF. This observation demonstrates the practical applicability of these metrics in scenarios where data influence evolves over time as a result of fine-tuning.
Deep Dive into RAG Noise Effect
We computed the mean similarities ($s_{S}$) and residuals ($r_{k,S}=s_{S\cup\{D_{k}\}}-s_{S}$) for the BoolQ datasets in all LLM, as shown in Table 8. These metrics provide a nuanced understanding of how data sets influence model outputs.
For the BoolQ dataset, the high negative residual for Bloomz (-0.32) indicates that adding the BoolQ context significantly influences the model output. This substantial change highlights the alignment of the data set with the pre-existing knowledge of the model, as seen in the high similarity score ($s_{S}=0.86$). In contrast, GPT-4’s low residual (-0.03) and high similarity score ($s_{S}=0.76$) suggest that it has been preexposed to similar data during training, resulting in minimal performance changes when BoolQ context is added. Mistral 7B, with a positive residual (0.06), demonstrates a strong reliance on the added BoolQ context, suggesting that it benefits greatly from this additional information. Similarly, Phi-3’s small positive residual (0.03) indicates partial exposure to similar data but with room for improvement when additional context is provided. We present the results on FakeQ and Olympics2024 data in Appendix E.
@@
Table 8: Mean similarities $s_{S}$ and residuals $r_{k,S}$ for different LLMs (with standard deviations).
LLM $s_{S}$ $s_{S\cup\{D_{k}\}}$ $r_{k,S}$
Bloomz 0.86 (0.03) 0.54 (0.02) -0.32 (0.02)
GPT-4 0.76 (0.02) 0.72 (0.02) -0.03 (0.02)
Mistral7B 0.73 (0.03) 0.69 (0.03) 0.06 (0.02)
Phi-3 0.60 (0.03) 0.63 (0.02) 0.03 (0.02)
@@ 5 Conclusion and Discussion
Our results show that both of our proposed algorithms successfully attribute the output of LLMs to the BoolQ dataset (as a proxy for related knowledge). Comparison of LLMs GPT-4 showed minimal change in similarities when the BoolQ context was added, suggesting prior exposure to similar data, while Bloomz exhibited a high residual, indicating substantial influence from BoolQ. The CMF algorithm provides further insight by quantifying the contributions of the base LLM. Comparing our two methods, we conclude that CMF is computationally less expensive and more robust to the RAG noise.
Our work still faces several limitations. First, our experiments were conducted on a few thousand queries, which allowed for controlled evaluation but may not fully capture the behavior of attribution metrics at scale. Scaling to significantly larger datasets—on the order of millions of queries—will introduce computational challenges, particularly in maintaining efficiency for CMF and SCM. We anticipate that while CMF’s structured decomposition will help mitigate noise at scale, the increased data volume may necessitate optimizations such as batched evaluations or adaptive sampling techniques.
Second, our evaluation was restricted to binary classification tasks, allowing for precise measurement of correctness and attribution effects. However, many real-world applications involve open-ended text generation, summarization, or multiclass classification. Extending our methods to these domains presents challenges in defining attribution signals, as generated outputs are more variable and harder to quantify. A key area of future work will be designing evaluation metrics that align dataset influence with generative model output, possibly leveraging perplexity shifts or token-level influence estimation.
Third, our RAG setup uses a 512-token chunk size, chosen based on the average passage length in BoolQ ( 108 tokens) to ensure coherent retrieval without excessive truncation or padding. While this is effective for our study, modern LLMs support much larger context sizes (e.g., 2500 to 300K tokens), which could improve retrieval for datasets with longer passages. Future work will explore how increasing chunk sizes impacts attribution stability, particularly in tasks requiring broader contextual understanding and long-form reasoning.
Appendix A Prompts
@@
Appendix B Olympic 2024 dataset
The Olympic 2024 dataset was constructed by sourcing textual snippets from publicly available Kaggle datasets Kaggle (2024) related to the Paris 2024 Olympics. Since no structured QA dataset existed for this event, question-answer pairs were manually generated for each context snippet to create a well-defined evaluation framework. For each retrieved context, three binary (Yes/No) question-answer pairs were created: one for training, one for validation, and one for testing. The training set was used to fine-tune the models, allowing them to incorporate Olympic-related information. The validation set was used to evaluate attribution metrics, ensuring that the methods assessed data influence without direct exposure during training. The test set was reserved for evaluating generalization after fine-tuning and unlearning experiments. This design follows the structure of BoolQ, making it easier to analyze how fine-tuning impacts attribution and how well unlearning methods reduce the model’s reliance on the dataset.
Appendix C Effectiveness of RAG
Table 9: Accuracy of LLMs with Different Contexts.
Context Bloomz GPT-4 Mistral 7B Phi-3
No Context 0.43 0.73 0.68 0.70
BoolQ as RAG 0.74 0.87 0.85 0.82
Five Datasets Only 0.35 0.64 0.60 0.45
All Data + BoolQ 0.73 0.84 0.82 0.83
The baseline setting (No Context) reveals inherent differences in the LLMs’ capabilities. GPT-4 has the highest baseline accuracy at 0.73, followed by Phi-3 at 0.70 and Mistral 7B at 0.68, indicating robust pretraining for these models. Bloomz shows lower accuracy at 0.43, highlighting its dependency on contextual data.
When the BoolQ context is provided using RAG, all models show significant accuracy improvements, with GPT-4 reaching 0.87, Mistral 7B at 0.85, and Phi-3 at 0.82. Bloomz also improves to 0.74, though it remains lower than the others. Providing context from five datasets (excluding BoolQ) leads to accuracy drops for all models, with Bloomz at 0.35, GPT-4 at 0.64, Mistral 7B at 0.60, and Phi-3 at 0.45. This indicates that less relevant data are less effective in understanding BoolQ queries.
Combining all data sets with the BoolQ context results in slight decreases in accuracy for GPT-4 (0.84) and Mistral 7B (0.82), suggesting that additional data sets introduce noise. Bloomz and Phi-3 show minimal changes, indicating that additional data do not significantly impact their performance once the BoolQ context is included. These results emphasize the importance of relevant contextual information in improving LLM performance, with GPT-4 consistently outperforming other models due to its extensive training.
Appendix D Additional Detailed Results
Tables 10 and 11 display the attribution results for the FakeQ dataset, which was constructed by altering the queries and contexts in BoolQ to ensure that it is unseen by the LLMs. SCM Attribution ($\phi_{\text{FakeQ}}$): The attribution values for FakeQ are noticeably lower than those for BoolQ. For example, $\phi_{\text{FakeQ}}$ ranges from 0.28 to 0.36 in LLM, compared to 0.48 to 0.59 for $\phi_{\text{BoolQ}}$. This reflects the lack of prior exposure to FakeQ in the training data, leading to a reduced alignment with the preexisting knowledge of the model.
CMF Attribution ($\pi_{\text{FakeQ}}$):
CMF assigns higher scores than SCM, with $\frac{\pi_{\text{FakeQ}}}{1-\pi_{\text{Base}}}$ ranging from 0.48 to 0.53 in LLM. This increase indicates the sensitivity of the CMF to the structural similarities between FakeQ and BoolQ. Although FakeQ is unseen, its construction retains the semantic patterns of BoolQ, allowing the models to leverage these structural similarities.
Tables 12 and 13 present the attribution results for the Olympic2024 dataset, which is entirely unseen and constructed to simulate real-world scenarios. The findings here are markedly different:
SCM Attribution ($\phi_{\text{Olympic2024}}$):
The Olympic2024 attribution values are significantly lower than both BoolQ and FakeQ. For example, $\phi_{\text{Olympic2024}}$ ranges from 0.07 to 0.11 in LLMs. This is expected since Olympic2024 is entirely unrelated to the models’ pre-training data, and its context does not align with the questions posed.
CMF Attribution ($\pi_{\text{Olympic2024}}$):
CMF similarly assigns lower attribution scores to Olympic2024 compared to BoolQ and FakeQ, with $\frac{\pi_{\text{Olympic2024}}}{1-\pi_{\text{Base}}}$ ranging from 0.15 to 0.20. However, the CMF values remain slightly higher than SCM, demonstrating its ability to account for small signal contributions even in unseen datasets. The behavior of the Olympic2024 attribution metrics highlights their reliability in distinguishing between datasets that are seen (BoolQ), partially similar (FakeQ), and entirely novel (Olympic2024). The low residuals for Olympic2024, particularly for models such as GPT-4, suggest minimal influence from prior training data, further validating the robustness of the attribution methods.
Metric Bloomz GPT-4 Mistral 7B Phi-3
$\phi_{\text{FakeQ}}$ 0.32 0.36 0.33 0.28
$\phi_{\text{Chemistry}}$ 0.08 0.06 0.07 0.08
$\phi_{\text{Natural Sci}}$ 0.09 0.07 0.08 0.09
$\phi_{\text{History}}$ 0.08 0.07 0.08 0.08
$\phi_{\text{Biology}}$ 0.07 0.05 0.06 0.07
$\phi_{\text{Law}}$ 0.06 0.05 0.06 0.06
Table 10: Shapley Values ($\phi_{k}$) using SCM Algorithm on FakeQ Dataset.
Table 11: $\bm{\pi}$ and $\frac{\pi_{\text{FakeQ}}}{1-\pi_{\text{Base}}}$ values using the CMF algorithm.
Metric Bloomz GPT-4 Mistral 7B Phi-3
$\pi_{\text{Base}}$ 0.06 0.09 0.07 0.06
$\pi_{\text{FakeQ}}$ 0.45 0.48 0.47 0.42
$\pi_{\text{Chemistry}}$ 0.10 0.07 0.08 0.09
$\pi_{\text{Natural Sci}}$ 0.08 0.07 0.07 0.08
$\pi_{\text{History}}$ 0.09 0.08 0.08 0.08
$\pi_{\text{Biology}}$ 0.07 0.06 0.06 0.06
$\pi_{\text{Law}}$ 0.06 0.06 0.06 0.06
$\frac{\pi_{\text{FakeQ}}}{1-\pi_{\text{Base}}}$ 0.49 0.53 0.51 0.48
Metric Bloomz GPT-4 Mistral 7B Phi-3
$\phi_{\text{Olympic2024}}$ 0.08 0.11 0.09 0.07
$\phi_{\text{Chemistry}}$ 0.04 0.03 0.04 0.05
$\phi_{\text{Natural Sci}}$ 0.05 0.04 0.04 0.05
$\phi_{\text{History}}$ 0.05 0.04 0.05 0.05
$\phi_{\text{Biology}}$ 0.04 0.03 0.03 0.04
$\phi_{\text{Law}}$ 0.03 0.02 0.03 0.03
Table 12: Shapley Values ($\phi_{k}$) using SCM Algorithm on Olympic2024 Dataset.
Table 13: $\bm{\pi}$ and $\frac{\pi_{\text{Olympic2024}}}{1-\pi_{\text{Base}}}$ values using the CMF algorithm.
Metric Bloomz GPT-4 Mistral 7B Phi-3
$\pi_{\text{Base}}$ 0.08 0.11 0.09 0.08
$\pi_{\text{Olympic2024}}$ 0.15 0.18 0.16 0.14
$\pi_{\text{Chemistry}}$ 0.07 0.05 0.06 0.06
$\pi_{\text{Natural Sci}}$ 0.06 0.05 0.06 0.06
$\pi_{\text{History}}$ 0.07 0.06 0.06 0.06
$\pi_{\text{Biology}}$ 0.06 0.05 0.05 0.05
$\pi_{\text{Law}}$ 0.05 0.05 0.05 0.05
$\frac{\pi_{\text{Olympic2024}}}{1-\pi_{\text{Base}}}$ 0.16 0.20 0.18 0.15
Appendix E Extended Deep Dive into RAG Noise Effect
For the FakeQ dataset, the results reveal its controlled construction and structural similarity to BoolQ. Bloomz exhibits a moderately negative residual (-0.16) and a slightly reduced similarity score ($s_{S}=0.78$), indicating that although FakeQ is not identical to BoolQ, its design allows the model to relate to it effectively. GPT-4 maintains a low residual (-0.03), reinforcing its robustness in handling data sets that resemble those encountered during training. Mistral 7B and Phi-3 show small positive residuals (0.03 and 0.04, respectively), suggesting that these models benefit from the added FakeQ context while exhibiting a less direct alignment compared to BoolQ.
For the Olympic2024 dataset, the results underscore its novelty. Bloomz has a less negative residual (-0.13) and a reduced similarity score ($s_{S}=0.72$), highlighting the limited alignment of this entirely novel data set with the model’s pre-existing knowledge. GPT-4 continues to display a low residual (-0.03), showing its robustness even when handling unseen data. Mistral 7B and Phi-3 exhibit small positive residuals (0.03 and 0.05, respectively), indicating their reliance on added context to improve performance. The lower similarity scores across all models for Olympic2024 reflect the unique nature of the data set, distinguishing it from the other data sets.
Table 14: Mean similarities $s_{S}$ and residuals $r_{k,S}$ for FakeQ across different LLMs (with standard deviations).
LLM $s_{S}$ $s_{S\cup\{D_{k}\}}$ $r_{k,S}$
Bloomz 0.78 (0.04) 0.62 (0.03) -0.16 (0.03)
GPT-4 0.69 (0.03) 0.66 (0.02) -0.03 (0.02)
Mistral 7B 0.65 (0.04) 0.68 (0.03) 0.03 (0.03)
Phi-3 0.58 (0.04) 0.62 (0.03) 0.04 (0.03)
Table 15: Mean similarities $s_{S}$ and residuals $r_{k,S}$ for Olympic2024 across different LLMs (with standard deviations).
LLM $s_{S}$ $s_{S\cup\{D_{k}\}}$ $r_{k,S}$
Bloomz 0.72 (0.05) 0.59 (0.04) -0.13 (0.04)
GPT-4 0.64 (0.03) 0.61 (0.03) -0.03 (0.03)
Mistral 7B 0.60 (0.04) 0.63 (0.03) 0.03 (0.03)
Phi-3 0.55 (0.05) 0.60 (0.04) 0.05 (0.04)